### v2

# Supervised heterogeneous network estimation via survival-based Bayesian graphical models

Abstract
Heterogeneity is a fundamental characteristic of cancer. To accommodate heterogeneity, subgroup identification has been extensively studied and broadly categorized into unsupervised and supervised analysis. Compared to unsupervised analysis, supervised approaches potentially hold greater clinical implications. Under the unsupervised analysis framework, several methods focusing on network-based subgroup identification have been developed, offering more comprehensive insights than those restricted to mean, variance, and other simplistic distributions by incorporating the interconnections among variables. However, research on supervised network-based subgroup identification remains limited. In this study, we develop a novel supervised Bayesian graphical model for jointly identifying multiple heterogeneous networks and subgroups. In the proposed model, heterogeneity is not only reflected in molecular data but also associated with a clinical outcome, and a novel similarity prior is introduced to effectively accommodate similarities among the networks of different subgroups, significantly facilitating clinically meaningful biological network construction and subgroup identification. The consistency properties of the estimates are rigorously established, and an efficient algorithm is developed. Extensive simulation studies and a real-world application to TCGA data are conducted, which demonstrate the advantages of the proposed approach in terms of both subgroup and network identification.
Keywords: Bayesian analysis; Finite mixture of regression models; Gaussian graphical model; Heterogeneous network estimation.
1 Introduction
For cancer, heterogeneity is a fundamental characteristic and is essential for unraveling the inherent complexities of biological systems. To describe and accommodate heterogeneity, researchers have conducted extensive studies on subgroup identification with high-dimensional molecular data, which can be roughly classified into unsupervised and supervised. Unsupervised approaches typically emphasize differences in the means of molecular factors (Table 1, a), with examples including [18] and [4], and variances (Table 1, b), as exemplified by [26]. Under the supervised analysis framework, a popular approach is the finite mixture of regression models, which captures the heterogeneity of genetic effects on clinical outcomes (Table 1, d). Examples include both Bayesian approaches [21, 1] and penalization approaches [24, 31]. These models enable the integration of information from both clinical outcomes and molecular data to identify clinically relevant subgroups. Despite significant achievements, existing techniques remain inadequate as they focus on the distributional properties or the influences of individual variables but neglect the interconnections among variables. There is a strong need for more informative approaches for subgroup identification.
Table 1: Comparison of unsupervised and supervised subgroup identification models.
Type Model and Example
Unsupervised a. Mean-based: $\bm{X}\sim\mathcal{N}(\bm{\mu}_{k},\bm{V})$, where $\bm{\mu}_{k}$ is the mean vector in the $k$th subgroup and $\bm{V}$ is the diagonal covariance matrix. Example: $\bm{\mu}_{1}=(1,0,0)^{\top},\ \bm{\mu}_{2}=(0,1,-1)^{\top}$.
b. Variance-based: $\bm{X}\sim\mathcal{N}(\bm{\mu}_{k},\bm{D}_{k})$, where $\bm{D}_{k}$ is the diagonal covariance matrix in the $k$th subgroup. Example: $\bm{D}_{1}=\text{diag}(1,2,1),\ \bm{D}_{2}=\text{diag}(2,1,2)$.
c. Network-based: $\bm{X}\sim\mathcal{N}(\bm{\mu}_{k},\bm{\Omega}_{k}^{-1})$, where $\bm{\Omega}_{k}$ is the precision matrix in the $k$th subgroup. Example: $\bm{\Omega}_{1}=\begin{pmatrix}1&-1&0\\ -1&2&0\\ 0&0&1\end{pmatrix},\ \bm{\Omega}_{2}=\begin{pmatrix}2&0&-1\\ 0&1&0\\ -1&0&2\end{pmatrix}$.
Supervised d. Linear effect-based: $Z\sim\mathcal{N}(\bm{\beta}_{k}^{\top}\bm{X},\sigma_{k}^{2})$, where $\bm{\beta}_{k}$ and $\sigma_{k}$ are the regression parameter vector and standard deviation in the $k$th subgroup. Example: $\bm{\beta}_{1}=(-1,0,0)^{\top},\ \bm{\beta}_{2}=(0,1,-1)^{\top};\ \sigma_{1}=1,\ \sigma_{2}=2$.
The importance of network-based subgroup identification has been emphasized in recent literature [22]. Subtype-specific networks are pivotal in understanding complex biological processes. They utilize information embedded in the interconnections among variables (Table 1, c), providing richer insights than marginal means and variances. As illustrated in the literature [7, 19], in addition to heterogeneity, there can be commonalities across subgroup-specific networks, which can help improve estimation efficiency and enhance interpretability through information borrowing.
Among the diverse statistical methods for network analysis, Gaussian graphical models (GGMs) have been prominent. Here, the nodes of a network are represented by random variables with a joint Gaussian distribution, while the edges reflect conditional dependencies between nodes that are determined by non-zero entries of the inverse covariance matrix (or precision matrix) [16, 10, 5]. To detect unknown heterogeneity, several network-based methods have been proposed under the framework of the Gaussian mixture model (GMM), with particular attention to identifying both structural differences and potential commonalities among multiple networks. For example, [11] incorporated a joint fused graphical lasso penalty into GMM for learning the related structures across different unknown subgroups. Similarly, a SCAN method proposed by [12] used a joint group graphical lasso penalty for simultaneous subgroup identification and joint graphical model estimation. [20] proposed decomposing the GGM into a set of sparse regression problems and introduced a composite minimax concave penalty for jointly estimating multiple networks. Under the Bayesian framework, [6] proposed a BCGS method that imposes a joint spike-and-slab graphical lasso prior on the precision matrices.
However, these GMM-based approaches belong to the unsupervised paradigm, focusing solely on the discovery of molecularly separable subgroups, do not consider any clinical outcomes, and may lead to subgroups lacking sufficient clinical implications. Published literature has demonstrated that supervised heterogeneity analysis that incorporates clinical information is promising in identifying clinically relevant subgroups and provides a more comprehensive understanding of disease [24, 31]. However, the existing supervised heterogeneity methods predominantly focus on detecting heterogeneous predictor-outcome associations without appropriately accommodating the network structures among predictors. Overall, there is a notable gap in jointly exploring molecular network structures and conducting subgroup identification in a supervised manner.
In this paper, we develop a novel supervised network-based heterogeneity analysis approach to simultaneously conduct network estimation and subgroup identification. This approach is among the first to conduct joint estimation for multiple networks in a supervised manner with unknown subgroup structures. Significantly advancing from the existing unsupervised network-based subgroup identification approaches [12, 14, 6], we develop a novel Bayesian framework integrating a finite mixture of Gaussian graphical models with supervised regression models to accommodate heterogeneity involving not only the distributions of predictors but also the predictor-outcome associations, leading to clinically more meaningful biological network construction and subgroup identification. Additionally, advancing from the existing joint network analysis based on the group penalties [7, 12, 14, 6], a novel similarity prior is introduced to capture the similarity information between the values of connection effects of distinct networks, facilitating more effective information borrowing and more appropriate modeling of the underlying biological mechanisms. Furthermore, unlike the existing supervised subgroup identification approaches [1, 31], we fully take the specific genetic networks into account while simultaneously accommodating the clinical information. The consistency properties of the proposed estimators are rigorously established, achieved through an efficient algorithm. Comprehensive simulation studies are conducted, demonstrating the advantages of the proposed approach in terms of subgroup and network identification. Additionally, a real-world application to TCGA data highlights the practical utilization of our approach.
2 Methods
We consider data with a survival outcome. Modeling survival presents more challenges due to non-negativity and censoring, and similar approaches can be applied to other clinical outcomes. Let $T$, $\Delta$, and $\bm{X}=\left(X_{1},\ldots,X_{p}\right)^{\top}$ represent the logarithm of the minimum of survival time and censoring time, censoring indicator, and $p$-dimensional vector of predictors (e.g., gene expressions), respectively. Here, $T=\operatorname{min}\left(Z,C\right)$ with $Z$ and $C$ being the logarithms of survival time and censoring time, and $\Delta={I}\{Z\leq C\}$ with ${I}\{\cdot\}$ being the indicator function. Assume there are $n$ subjects $\left\{\left(t_{i},\delta_{i},\bm{x}_{i}\right),i=1,\ldots,n\right\}$, which form $K$ subgroups. Denote $\bm{z},\bm{c},\bm{\delta},$ and $\bm{t}$ as the vectors of $z_{i}$’s, $c_{i}$’s, $\delta_{i}$’s, and $t_{i}$’s, respectively, and X as the $n\times p$ matrix composed of $\bm{x}_{i}$’s.
2.1 Bayesian mixture modeling
Denote $g_{i}\in\{1,\ldots,K\}$ as the subgroup membership of the $i$th subject. Consider the accelerated failure time (AFT) model:
[…]
where $\ddot{\beta}_{0k}$ and $\ddot{\bm{\beta}}_{k}=\left(\ddot{\beta}_{k,1},\ldots,\ddot{\beta}_{k,p}\right)^{\top}$ are the subgroup-specific intercept and regression parameters, and $\epsilon_{i}$ is a normally distributed random variable with mean 00 and precision parameter ${\tau}_{k}^{2}$. We propose the following Bayesian mixture model with a reparameterization technique:
$t_{i}\mid\left(\delta_{i},z_{i},c_{i}\right)=z_{i}^{\delta_{i}}c_{i}^{1-\delta_{i}},\delta_{i}\mid\left(z_{i},c_{i}\right)={I}\left\{z_{i}\leq c_{i}\right\},c_{i}\sim f_{C}(c_{i}),$ (1)
$z_{i}\mid\left(\bm{\beta},\bm{x}_{i},g_{i}=k;\bm{\tau}\right)\sim\mathcal{N}\left(z_{i}\mid\beta_{0k}/\tau_{k}+\bm{\beta}_{k}^{\top}\bm{x}_{i}/\tau_{k},\tau_{k}^{-2}\right),\bm{x}_{i}\mid\left(\bm{\mu},\bm{\Omega},g_{i}=k\right)\sim\mathcal{N}\left(\bm{x}_{i}\mid\bm{\mu}_{k},\bm{\Omega}_{k}^{-1}\right).$
Here, $f_{C}$ is the density function of the logarithm of censoring time $C$, $\bm{\mu}_{k}=\left(\mu_{k,1},\cdots,\mu_{k,p}\right)^{\top}$ and $\bm{\Omega}_{k}=\left(\omega_{k,jl}\right)_{p\times p}$ are the mean and precision matrix for the predictors in the $k$th subgroup, and $\bm{\beta}$, $\bm{\tau}$, $\bm{\mu}$, and $\bm{\Omega}$ are the vectors of $\left({\beta}_{0k},\bm{\beta}_{k}\right)$’s, ${\tau}_{k}$’s, $\bm{\mu}_{k}$’s, and $\bm{\Omega}_{k}$’s, respectively, with ${\bm{\beta}}_{k}=\ddot{\bm{\beta}}_{k}\tau_{k}$ and $\beta_{0k}=\ddot{\beta}_{0k}\tau_{k}$. For the $k$th subgroup, we estimate the network based on $\bm{\Omega}_{k}$, where $\omega_{k,jl}\neq 0$ corresponds to a connection between predictors $l$ and $j$.
After integrating out $\bm{z}$ and $\bm{c}$, we have the joint distribution of $\left(\bm{t},\bm{\delta},\textbf{X}\right)$ as:
$f\left(\bm{t},\bm{\delta},\textbf{X}\mid\bm{\beta},\bm{\mu},\bm{\Omega},\bm{\tau},\bm{g}\right)=\prod_{i=1}^{n}\prod_{k=1}^{K}\left[f_{k}\left(t_{i},\bm{x}_{i},\delta_{i}\mid{\bm{\mu}}_{k},{\bm{\Omega}}_{k},{\bm{\beta}}_{k},{\beta}_{0k},{{\tau}}_{k}\right)\right]^{I\left\{g_{i}=k\right\}}$
$=$ $\prod_{i=1}^{n}\prod_{k=1}^{K}\left\{f_{\bm{X}}\left(\bm{x}_{i}\mid\bm{\mu}_{k},\bm{\Omega}_{k}\right)\left[f_{Z}\left(t_{i}\mid\bm{x}_{i},\bm{\beta}_{k},\beta_{0k},{\tau}_{k}\right)S_{C}(t_{i})\right]^{\delta_{i}}\left[f_{C}(t_{i})S_{Z}\left(t_{i}\mid\bm{x}_{i},\bm{\beta}_{k},\beta_{0k},{\tau}_{k}\right)\right]^{1-\delta_{i}}\right\}^{{I}\left\{g_{i}=k\right\}},$
where $f_{\bm{X}}$ and $f_{Z}$ are the density functions of $\bm{X}$ and $Z$ given in (1), respectively, and $S_{C}$ and $S_{Z}$ are the survival functions of $C$ and $Z$, respectively.
In (1), we assume random censoring with $c_{i}$ being independent of $z_{i}$ and $\bm{x}_{i}$, as in the literature [24]. We propose combining the AFT model with the mixture Gaussian graphical model to accommodate the multi-layer heterogeneity. Specifically, in the AFT model, we introduce a subgroup-specific coefficient $\bm{\beta}_{k}$ to explicitly capture the heterogeneous associations of the molecular predictors with survival time. Here, a reparameterization technique based on the noise level is adopted for $\bm{\beta}_{k}$ to ensure that the proposed estimator remains scale-invariant under the mixture model [25, 31]. This technique facilitates an “adaptive adjustment” of the coefficients, aligning them with data characteristics and leading to more accurate estimations. On the other hand, within the Gaussian graphical model, we utilize $\bm{\mu}_{k}$ and $\bm{\Omega}_{k}$ to characterize heterogeneity attributed to both differential mean expressions and differential network structures. Consequently, the network’s heterogeneity is not only intricately linked to the interested measurements but also closely connected to the distribution of survival time, making the resulting networks clinically more meaningful. For example, researchers can better understand the differences between patients with good survival and those with bad survival on specific biological networks, thereby more effectively guiding subsequent treatments. Here, to simplify our approach, consistent with prior studies [12, 8], we assume that the subgroup number $K$ is identifiable.
2.2 Prior specification
We propose the following priors for the parameters in (1):
${\omega_{k,jj}\sim\operatorname{Exp}(\omega_{k,jj};\tau_{0})},\text{ for }j=1,\ldots,p,$
$\omega_{k,jl}\mid\gamma_{k,jl}\sim\gamma_{k,jl}\operatorname{LP}\left(\omega_{k,jl};v_{1}\right)+\left(1-\gamma_{k,jl}\right)\operatorname{LP}\left(\omega_{k,jl};v_{0}\right),\gamma_{k,jl}\sim\operatorname{Bern}\left(p_{1}\right),\text{ for }j<l,$
$\beta_{k,j}\sim\operatorname{LP}\left(\beta_{k,j};1/\lambda_{1}\right),\mu_{k,j}\sim\operatorname{LP}\left(\mu_{k,j};1/\lambda_{2}\right),g_{i}\mid\bm{\pi_{g}}\sim\operatorname{Multi}\left(\pi_{1},\pi_{2},\ldots,\pi_{K}\right).$
Here, $\operatorname{Exp}(\omega_{k,jj};\tau_{0})$ is the Exponential distribution with parameter $\tau_{0}$, $\gamma_{k,jl}$ is a binary indicator corresponding to $\omega_{k,jl}$, $\operatorname{LP}(y;v)$ is the Laplace distribution with a scale parameter $v$, $p_{1}\in(0,1)$ is the prior probability for the Bernoulli distribution, and $\bm{\pi}_{g}=\left(\pi_{1},\ldots,\pi_{K}\right)^{\top}$ is the vector composed of the Multinomial distribution’s parameters, which specifies the prior probabilities that each sample is drawn from different subgroups. $\tau_{0}$, $\lambda_{1}$, $\lambda_{2}$, and $v_{1}>v_{0}$ are the positive tuning parameters.
In the proposed Bayesian framework, we first introduce a weakly informative Exponential prior to the diagonal elements $\omega_{k,jj}$’s of the precision matrices, which are not subject to shrinkage, following [10]. We introduce a spike-and-slab Laplace prior [23, 28] for $\omega_{k,jl}$ based on the indicator $\gamma_{k,jl}$, for identifying individual sparse patterns of the $K$ networks. With this prior, when $\gamma_{k,jl}=1$, the corresponding entry $\omega_{k,jl}$ comes from the slab component with a relative large variance $2v_{1}^{2}$ and is likely to take a value away from zero; when $\gamma_{k,jl}=0$, $\omega_{k,jl}$ comes from the spike component with a small variance $2v_{0}^{2}$ that can induce values towards zero. Compared to the Laplace prior, the spike-and-slab Laplace prior involves two different scale parameters $v_{1}$ and $v_{0}$ to distinguish active from ignorable coefficients and can nonlinearly adapt to coefficients with $\gamma_{k,jl}$, achieving “selective shrinkage” with a smaller bias. This Laplace distribution-based prior can automatically achieve sparsity via the $L_{1}$ norm on $\omega_{k,jl}$. This advances from the commonly adopted spike-and-slab Gaussian prior which yields nonsparse posterior modes for $\omega_{k,jl}$’s and relies on a post-data thresholding-based selection strategy utilizing $\gamma_{k,jl}$. The strategy that is based on the estimation and sparsity of the continuous $\omega_{k,jl}$ can lead to more effective identification of sparse posterior patterns [23] and can also facilitate a simpler theoretical study and a more efficient computation.
Additionally, to accommodate the sparse structure of the truly associated predictors and predictor means, sparse priors are assumed for ${\beta}_{k,j}$’s and $\mu_{k,j}$’s for regularized estimation. As the estimation of ${\beta}_{k,j}$’s and $\mu_{k,j}$’s is less challenging than $\omega_{k,jl}$, we choose the relatively simple Laplace prior for both the regression coefficients and predictor means, which involves fewer parameters and less computational complexity.
2.3 Additional priors for accommodating common structures
To borrow information from cross-subgroup similarity, we further introduce a latent vector $\bm{\theta}_{jl}$ for $\bm{\omega}_{jl}$ and assume a Gaussian distribution with the precision matrix based on a similarity-based matrix $\bm{L}^{(jl)}$,
$\bm{\theta}_{jl}\sim\mathcal{N}\left(\bm{0},\left(u\bm{L}^{(jl)}\right)^{-1}\right),\bm{\theta}_{jl}\mid\bm{\omega}_{jl}\sim{I}\left\{\bm{\theta}_{jl}=\bm{\omega}_{jl}\right\},$
with $u$ being the positive tuning parameter. Here, $\bm{\theta}_{jl}\sim\mathcal{N}\left(\bm{0},\left(u\bm{L}^{(jl)}\right)^{-1}\right)$ can be interpreted as a generative model in which a zero vector is observed from a Gaussian distribution parameterized by $\bm{\theta}_{jl}$, i.e., $f\left(\bm{0}\mid\bm{\theta}_{jl}\right)=\sqrt{u\left|\bm{L}^{(jl)}\right|/(2\pi)^{K}}\exp\left\{-u\left(\bm{0}-\bm{\theta}_{jl}\right)^{\top}\bm{L}^{(jl)}\left(\bm{0}-\bm{\theta}_{jl}\right)/2\right\}$. This interpretation follows the hierarchical framework of [30], where priors are constructed through pseudo-observations. We define $L^{(jl)}_{k,k}=(K-1)/\left(\theta_{k,jl}^{2}+\epsilon^{2}\right)$ and $L^{(jl)}_{k,k^{\prime}}=-1/\sqrt{\left(\theta_{k,jl}^{2}+\epsilon^{2}\right)\left(\theta_{k^{\prime},jl}^{2}+\epsilon^{2}\right)}$ for $1\leq k\neq k^{\prime}\leq K$ and $1\leq j<l\leq p$, with $\epsilon$ being a small constant. Thus, for each $\bm{\theta}_{jl}=(\theta_{1,jl},\ldots,\theta_{K,jl})^{\top}$, the prior involves $\bm{\theta}_{jl}^{\top}\bm{L}^{(jl)}\bm{\theta}_{jl}=0.5\sum_{k\neq k^{\prime}}\left(\theta_{k,jl}/\sqrt{\theta_{k,jl}^{2}+\epsilon^{2}}-\theta_{k^{\prime},jl}/\sqrt{\theta_{k^{\prime},jl}^{2}+\epsilon^{2}}\right)^{2}$, which is an approximation to the Bayesian counterpart of the sign-based constraint penalty, that is $0.5\sum_{k\neq k^{\prime}}\left[\operatorname{sgn}(\theta_{k,jl})-\operatorname{sgn}(\theta_{k^{\prime},jl})\right]^{2}$. Combined with the indicator function ${I}\left\{\bm{\theta}_{jl}=\bm{\omega}_{jl}\right\}$, the proposed approach encourages the pairwise relationships of predictors across subgroups to have similar signs, leading to networks with a shared pattern of sparsity as well as edges with shared positive or negative relationships. This significantly advances from the previous research based on the group penalty [7, 12, 14, 6], which utilizes group selection without appropriately accounting for the similarity of the estimated values and fails to sufficiently address the relationships between different networks [15]. This is also more flexible than the strategy based on fused lasso [7, 11, 29], which promotes both shared directions and magnitudes of edges. Similar prior assumptions can also be applied to $\bm{\beta}$ and $\bm{\mu}$ when considering the common structures within regression coefficients and predictors’ means.
2.4 Posterior Inference
The complete posterior distribution of the parameters conditionally on the available data is:
$f\left(\bm{\beta},\bm{\mu},\bm{\Omega},\bm{\theta},\bm{\gamma},\bm{g}\mid\bm{t},\bm{\delta},\textbf{X};\bm{\tau},\bm{\pi_{g}}\right)\propto$ $f\left(\bm{t},\bm{\delta},\textbf{X}\mid\bm{\beta},\bm{\mu},\bm{\Omega},\bm{\tau},\bm{g}\right)f\left(\bm{\beta}\right)f\left(\bm{\mu}\right)f\left(\bm{\Omega}\mid\bm{\gamma}\right)$ (2)
×$\times f\left(\bm{\gamma}\right)f\left(\bm{\theta}\mid\bm{\Omega}\right)f\left(\bm{0}\mid\bm{\theta}\right)f\left(\bm{g}\mid\bm{\pi_{g}}\right),$
with $f\left(\bm{0}\mid\bm{\theta}\right)=\prod_{j=1}^{p}\prod_{l>j}f\left(\bm{0}\mid\bm{\theta}_{jl}\right)$. We note that since the distribution $f\left(\bm{\theta}\mid\bm{\Omega}\right)$ is the indicator function, there is no need to estimate $\bm{\theta}$ separately. Figure S1 in the supplementary materials gives a graphical representation of the Bayesian model.
Based on (2), we develop an Expectation Maximization (EM) algorithm for posterior inference of the parameter set $\bm{\Psi}=\operatorname{vec}\left(\bm{\beta},\bm{\tau},\bm{\mu},\bm{\Omega}\right)$. Sensitivity analyses on hyperparameters confirm that our approach is robust across a range of plausible settings. The Bayesian Information Criterion (BIC) is adopted to choose optimal values of $v_{0}$ and $u$. We refer to Sections S2 and S3 of the supplementary materials for details.
2.5 Theoretical properties
Denote $\bm{\Psi}^{*}=\operatorname{vec}\left(\bm{\beta}^{*},\bm{\tau}^{*},\bm{\mu}^{*},\bm{\Omega}^{*}\right)$ as the vector composed of the unknown true parameters. For the $k$th subgroup, denote $\mathcal{U}_{k}^{*}=\left\{j:\mu_{k,j}^{\ast}\neq 0\right\}$ and $\mathcal{V}_{k}^{*}=\left\{j:\beta_{k,j}^{\ast}\neq 0\right\}$ as the sets of the nonzero elements of the true predictor means and regression parameters, and $\mathcal{W}_{k}^{*}=\left\{(j,l):j\neq l,\omega_{k,jl}^{\ast}\neq 0\right\}$ as the set of the nonzero off-diagonal elements of the true precision matrix. For a set $\mathcal{S}$, denote $\left|\mathcal{S}\right|$ as its cardinality. Define the sparsity parameters as $d_{\bm{\mu}}=\left|\cup_{k=1}^{K}\mathcal{U}_{k}^{*}\right|,d_{\bm{\beta}}=\left|\cup_{k=1}^{K}\mathcal{V}_{k}^{*}\right|$, and $s=\left|\cup_{k=1}^{K}\mathcal{W}_{k}^{*}\right|$. Define $\Xi$ as some non-empty convex set of parameters $\bm{\Psi}$ in $\mathbb{R}^{Kp^{2}+2Kp+2K}$. Denote ${\rho}_{k,i}^{\bm{\Psi}}$ as the probability that the $i$th subject belongs to the $k$th subgroup given $\bm{{\Psi}}$.
Furthermore, for a $q$-dimensional vector $\bm{\nu}$, denote $\left\|\bm{\nu}\right\|_{\max}=\max_{1\leq j\leq q}\left|\nu_{j}\right|$. For a matrix $\bm{M}\in\mathbb{R}^{q_{1}\times q_{2}}$, define its minimum, maximum, and max-induced norms as $\|\bm{M}\|_{\min}=\min_{i,j}\left|M_{ij}\right|$, $\|\bm{M}\|_{\max}=\max_{i,j}\left|M_{ij}\right|$, and $\|\bm{M}\|_{\infty}=\max_{i=1,\ldots,q_{1}}\sum_{j=1}^{q_{2}}\left|M_{ij}\right|$, respectively. For a square matrix $\bm{A}\in\mathbb{R}^{p\times p}$, let $\lambda_{\min}(\bm{A})$ and $\lambda_{\max}(\bm{A})$ denote its smallest and largest eigenvalues. The following conditions are assumed.
Assumption 1
Suppose $\underline{C}_{1}\leq\min_{k}\lambda_{\min}\left(\bm{\Omega}_{k}^{*}\right)\leq\max_{k}\lambda_{\max}\left(\bm{\Omega}_{k}^{*}\right)\leq\bar{C}_{1}$, where $\underline{C}_{1}$ and $\bar{C}_{1}$ are positive constants. Moreover, we denote ¨$\ddot{\ell}_{1}\left(\bm{\Psi},\bar{\bm{\Psi}}\right)$ and ¨$\ddot{\ell}_{2}\left(\bm{\Psi},\bar{\bm{\Psi}}\right)$ as the second derivatives of $\sum_{i}\sum_{k}{\rho}_{k,i}^{\bm{\Psi}}\delta_{i}\log f_{Z}\left(t_{i}\mid\bm{x}_{i},\bar{\bm{\beta}}_{k};\bar{\beta}_{0k},\bar{\tau}_{k}\right)$ and $\sum_{i}\sum_{k}{\rho}_{k,i}^{\bm{\Psi}}(1-\delta_{i})\log S_{Z}\left(t_{i}\mid\bm{x}_{i},\bar{\bm{\beta}}_{k};\bar{\beta}_{0k},\bar{\tau}_{k}\right)$ with respect to $\operatorname{vec}\left({\bm{\beta}},\bm{\tau}\right)$, respectively. Then for any $\underline{\bm{\Psi}}\in\Xi$, $\bm{\Psi}\in\mathcal{B}_{\alpha}\left(\bm{\Psi}^{*}\right)=\left\{\bm{\Psi}\in\Xi:\left\|\bm{\Psi}-\bm{\Psi}^{*}\right\|_{2}\right.$ $\left.\leq\alpha\right\}$ and $\bar{\bm{\Psi}}\in\left\{\bar{\bm{\Psi}}\mid\left\|\bar{\bm{\Psi}}-\bm{\Psi}^{*}\right\|_{2}\leq 2\alpha\right\}$, there exists $\widetilde{\eta}>0$ satisfying $\lambda_{\max}\left(\ddot{\ell}_{1}\left(\bm{\Psi},\bar{\bm{\Psi}}\right)\right)\leq-\widetilde{\eta}$ and $\lambda_{\max}\left(\ddot{\ell}_{2}\left(\bm{\Psi},\underline{\bm{\Psi}}\right)\right)\leq 0.$
Assumption 2
[…]
Assumption 3
$\left\|\bm{\beta}^{*}\right\|_{\max}$, $\left\|\bm{\tau}^{\ast}\right\|_{\max}$,
$\left\|\bm{\mu}^{*}\right\|_{\max}$, and $\max_{k}\left\|\bm{\Omega}_{k}^{*}\right\|_{\infty}$ are bounded from above, and $\left\|\bm{\tau}^{\ast}\right\|_{\min}$ is bounded from below. Furthermore, $\max\limits_{i=1,\cdots,n}{c}_{i}\leq\bar{C}_{2}$ and $\max\limits_{i=1,\cdots,n}\left\|\bm{x}_{i}\right\|_{\max}\leq\bar{C}_{3}$ for some constants $\bar{C}_{2}$ and $\bar{C}_{3}$.
Assumption 4
The number of subgroups $K$ satisfies $K\sqrt{\log n/p}=o\left(1\right)$.
Assumption 5
Assume the tuning parameters satisfy $u=O\left(\sqrt{{n(Ks+p)\log p}/{(Kp^{2}s^{2})}}\right)$, $\tau_{0}=O\left(\sqrt{n{(Ks+p)K\log p}/p}\right)$, $1/v_{0}=O\left(\sqrt{nK(Ks+p)\log p/s}\right)$, $1/v_{1}=O\left(\sqrt{nK(Ks+p)\log p/s}\right)$, and $v_{0}<v_{1}$.
Assumption 1 is a restricted strong concave condition for the log-likelihood function [13, 2]. Assumption 2 is the sufficient separability condition, which requires that the probability of each subject belonging to a subgroup is either close to 1 or 0. Such an assumption is often assumed in the literature [12, 22]. Assumption 3 puts constraints on the bounds of the true parameters and the predictors, which are very common in the relevant literature. Assumption 4 allows $K$ to grow at a rate slower than $\sqrt{p/\log n}$, which is also considered in the related literature [17, 8]. Assumption 5 restricts the rate of the tuning parameters, which are commonly assumed in high-dimensional studies [28, 31].
Theorem 1
Suppose that Assumptions 1-5 hold. Let $\widetilde{s}=d_{\bm{{\beta}}}\sqrt{d_{\bm{{\mu}}}}+d_{\bm{{\beta}}}^{2}+\sqrt{p}+\sqrt{{s+p/K}}$. Assume that $\widetilde{s}\sqrt{{K^{3}\log p}}=o(\sqrt{n})$ and the lower bound of the minimal signal in the true parameters satisfies $\min\limits_{k=1,\ldots,K}\left\{\min\limits_{(j,l)\in\mathcal{W}_{k}^{*}}|\omega_{k,jl}^{*}|,\min\limits_{j\in\mathcal{U}_{k}^{*}}|\mu_{k,j}^{*}|,\min\limits_{j\in\mathcal{V}_{k}^{*}}|\beta_{k,j}^{*}|\right\}>2\widetilde{s}\sqrt{{K^{3}\log p}/{n}}$. Then,
1. The non-asymptotic bound of the estimation error of $\hat{\bm{{\Psi}}}$ obtained with the proposed EM algorithm is
$\left\|\hat{\bm{\Psi}}-\bm{\Psi}^{*}\right\|_{2}=O_{p}\left(\widetilde{s}\sqrt{{K^{3}\log p}/{n}}\right).$ (3)
2. For the final estimators $\widetilde{\bm{\Omega}}_{k}$, $\widetilde{\bm{\mu}}_{k}$, and $\widetilde{\bm{\beta}}_{k}$ with $\widetilde{\omega}_{k,jl}=\hat{\omega}_{k,jl}{I}\left\{\left|\hat{\omega}_{k,jl}\right|>\widetilde{s}\sqrt{{K^{3}\log p}/{n}}\right\}$, $\widetilde{\mu}_{k,j}=\hat{\mu}_{k,j}{I}\left\{\left|\hat{\mu}_{k,j}\right|>\widetilde{s}\sqrt{{K^{3}\log p}/{n}}\right\}$, and $\widetilde{\beta}_{k,j}=\hat{\beta}_{k,j}{I}\left\{\left|\hat{\beta}_{k,j}\right|>\widetilde{s}\sqrt{{K^{3}\log p}/{n}}\right\}$, we have $\left\|\widetilde{\bm{\Omega}}-\bm{\Omega}^{*}\right\|_{2}=O_{p}\left(\widetilde{s}\sqrt{{K^{3}\log p}/{n}}\right)$ and with probability approaching 1, $\widehat{\mathcal{W}}_{k}=\mathcal{W}_{k}^{*}$, $\widehat{\mathcal{U}}_{k}=\mathcal{U}_{k}^{*}$, and $\widehat{\mathcal{V}}_{k}=\mathcal{V}_{k}^{*}$ for any $k=$ $1,\ldots,K$, where $\widetilde{\bm{\Omega}}=\operatorname{vec}\left(\widetilde{\bm{\Omega}}_{1},\ldots,\widetilde{\bm{\Omega}}_{K}\right)$, and $\widehat{\mathcal{W}}_{k}$ denotes the set of the nonzero off-diagonal elements of $\widetilde{\bm{\Omega}}_{k}$, and $\widehat{\mathcal{U}}_{k}$ and $\widehat{\mathcal{V}}_{k}$ denote the sets of the nonzero elements of $\widetilde{\bm{\mu}}_{k}$ and $\widetilde{\bm{\beta}}_{k}$, respectively.
In Theorem 1, if $K$ is fixed, the non-asymptotic bound of the estimation error reduces to $\left\|\widetilde{\bm{\Omega}}-\bm{\Omega}^{*}\right\|_{2}=O_{p}\left(\left(d_{\bm{{\beta}}}\sqrt{d_{\bm{{\mu}}}}+d_{\bm{{\beta}}}^{2}+\sqrt{{s+p}}\right)\sqrt{\log p/n}\right)$. Denote by $\mathcal{\bm{Q}}\equiv\left\{\bm{\Omega}=\operatorname{vec}\left(\bm{\Omega}_{1},\ldots,\right.\right.$ $\left.\left.\bm{\Omega}_{K}\right):\right.$ $\left.\bm{\Omega}_{k}\text{ is positive definite and~{}}\max_{k}\|\bm{\Omega}_{k}\|_{\infty}\leq C_{\mathcal{Q}}\right\}$ the class of precision matrices. Under Assumption 3, the upper bound $C_{\mathcal{Q}}$ of $\mathcal{\bm{Q}}$ is independent of $(n,p)$. Provided that $d_{\bm{{\beta}}}\sqrt{d_{\bm{{\mu}}}}+d_{\bm{{\beta}}}^{2}=O(\sqrt{s+p})$, the convergence rate of the precision matrix estimate $\left\|\widetilde{\bm{\Omega}}-\bm{\Omega}^{*}\right\|_{2}=O_{p}(\sqrt{(s+p)\log p/n})$, which achieves the minimax optimal rate for estimating an $s$-sparse precision matrix under the Frobenius norm as shown in [3]. The proof of Theorem 1 is provided in Section S5 of the supplementary materials.
3 Simulation
Simulation studies are conducted to evaluate the performance of the proposed approach under the following settings. (a) $K=2$ and $3$ and $p=100$. (b) Consider two settings for the sample size. One is a balanced design where there are 150 subjects in each subgroup, while the other is an imbalanced design with sample sizes being $100$ and $200$ for $K=2$ and $100,150$, and $200$ for $K=3$. Thus, there are $n=300$ subjects for $K=2$, and $n=450$ for $K=3$ in total. (c) The predictor observations $\left\{\bm{x}_{i}\right\}$ in the $k$th subgroup are simulated from $\mathcal{N}\left(\bm{0},\bm{\Omega}_{k}^{-1}\right)$, where $\bm{\Omega}_{k}$ is generated based on the corresponding network structure. Specifically, following [7], we simulate each network with ten unconnected subnetworks. To comprehensively examine the performance of the proposed approach under different degrees of similarity across subgroup-specific precision matrices, we consider three specific settings S1-S3, with the number of subnetworks sharing the same sparsity structure in all subgroups being 3, 5, and 7, respectively. For each subnetwork, we consider three types of structure, including the power-law, nearest-neighbor, and Erdös-Rényi networks. The detailed settings are provided in Section S6 of the supplementary materials. (d) Each logarithmic survival observation $t_{i}$ is generated as follows. First, the logarithmic survival times of the $k$th subgroup are generated from $\mathcal{N}\left(\bm{\beta}_{k}^{\top}\bm{x}_{i},0.01^{2}\right)$, where five elements of $\bm{\beta}_{k}$ are non-zero, and the rest $p-5$ elements are all zero with $\bm{\beta}_{1}=\left(2,2,2,2,2,0,\ldots,0\right)^{\top}$, $\bm{\beta}_{2}=-\left(2,2,2,2,2,0,\ldots,0\right)^{\top}$, and $\bm{\beta}_{3}=\left(0,0,1,1,1,1,1,0,\ldots,0\right)^{\top}$, respectively. Then, the censoring time is simulated from a Gamma distribution with the shape and scale parameters controlling the censoring rate to be $\sim 20\%$. There are a total of 36 scenarios.
In addition to the proposed approach, six alternative approaches are also considered. BCGS is a Bayesian approach for simultaneously clustering and learning multiple sparse graphical models using a joint spike-and-slab graphical lasso prior [6]. SCAN is a mixture Gaussian graphical model-based approach that applies the group lasso penalty to accommodate common structures across unknown heterogeneous subgroups [12]. TGemBag and TJGL are two “ideal approaches” which assume that the true subgroup memberships are known and adopt GemBag and JGL for jointly estimating multiple networks for the true subgroups, respectively. Here, GemBag is a Bayesian graphical model for accommodating similarity as well as difference of multiple networks [28], and JGL employs a likelihood-based joint graphical Lasso for multiple network estimations [7]. FMGemBag and FMJGL are two two-stage approaches, where in the first stage, a finite mixture (FM) of AFT models developed in [24] is used to obtain the subgroup memberships of all subjects, which can account for right censoring and unknown heterogeneous predictor-survival associations, and in the second stage, GemBag and JGL are adopted for network estimation based on the estimated subgroup memberships, respectively. Among these methods, BCGS, SCAN, TJGL, and FMJGL can be realized using or extending the functionalities provided by the R package JGL, and TGemBag and FMGemBag can be realized using R codes at https://github.com/xinming104/GemBag. BCGS and SCAN achieve identification of unknown heterogeneity and network structures simultaneously, however, are unsupervised and only account for predictors. GemBag and JGL are developed for estimating precision matrices with known subgroup memberships. We consider two types of subgroup memberships for GemBag and JGL, including the one based on the FM approach, which utilizes the information of survival for heterogeneous analysis, and the “ideal” type based on the true subgroup memberships.
To evaluate the performance of different approaches, we consider the clustering error (CE) for evaluating the heterogeneity identification performance, precision matrix squared error (PME) for evaluating the estimation performance, and true and false positive rates (TPR and FPR) for evaluating the network identification performance. The details are provided in Section S6 of the supplementary materials. Here, similar to other network-based heterogeneity analysis studies [12], we focus on the performance of estimation and identification on $\omega_{k,jl}$’s, which can be more challenging than on $\mu_{k,j}$’s and $\beta_{k,j}$’s.
100 replicates are conducted for each scenario, where the true value of $K$ is used for all approaches. The summary results for $K=2$ under power-law and Erdös-Rényi networks are presented in Tables 2-3, and the rest of the results are provided in Section S6 of the supplementary materials. Under all simulated scenarios, the proposed approach consistently demonstrates superior or competitive performance compared to the alternative approaches. In terms of subgroup identification, the proposed approach performs significantly better with much smaller CEs. Among the alternatives, the FM methods tend to exhibit the poorest subgroup identification performance, while SCAN and BCGS have better subgroup identification accuracy but are still much inferior compared to the proposed approach.
Table 2: Simulation results for the power-law network under the scenarios with $K=2$. In each cell, mean (SD) based on 100 replicates.
proposed BCGS SCAN TGemBag TJGL FMGemBag FMJGL
S1 with the balanced design
CE 0.001(0.003) 0.044(0.134) 0.045(0.134) - - 0.196(0.116) 0.196(0.116)
PME 4.502(0.362) 5.060(1.793) 5.054(1.784) 3.870(0.507) 7.789(0.528) 7.690(2.508) 8.695(1.577)
TPR 0.932(0.039) 0.901(0.142) 0.897(0.142) 0.933(0.043) 0.931(0.048) 0.871(0.128) 0.887(0.098)
FPR 0.025(0.005) 0.032(0.008) 0.031(0.008) 0.024(0.006) 0.025(0.008) 0.122(0.018) 0.080(0.022)
S2 with the balanced design
CE 0.002(0.005) 0.087(0.169) 0.088(0.170) - - 0.263(0.119) 0.263(0.119)
PME 4.398(0.335) 6.847(10.771) 5.453(2.296) 3.850(0.539) 7.821(0.583) 7.609(1.913) 8.581(1.299)
TPR 0.936(0.039) 0.870(0.174) 0.863(0.178) 0.932(0.051) 0.924(0.056) 0.866(0.116) 0.889(0.079)
FPR 0.025(0.006) 0.031(0.009) 0.030(0.008) 0.023(0.006) 0.024(0.009) 0.117(0.028) 0.081(0.032)
S3 with the balanced design
CE 0.009(0.010) 0.190(0.226) 0.195(0.230) - - 0.293(0.132) 0.293(0.132)
PME 4.134(0.348) 7.335(4.777) 6.917(3.424) 4.479(1.614) 8.222(1.080) 6.343(1.715) 8.367(3.251)
TPR 0.951(0.027) 0.775(0.255) 0.779(0.232) 0.867(0.172) 0.894(0.097) 0.910(0.123) 0.931(0.070)
FPR 0.023(0.006) 0.033(0.010) 0.033(0.009) 0.021(0.009) 0.022(0.011) 0.114(0.019) 0.074(0.022)
S1 with the imbalanced design
CE 0.001(0.002) 0.015(0.042) 0.015(0.039) - - 0.248(0.047) 0.248(0.047)
PME 4.929(0.446) 5.131(0.684) 5.147(0.685) 4.299(0.630) 7.806(0.514) 8.862(2.024) 9.379(1.350)
TPR 0.897(0.047) 0.877(0.084) 0.870(0.085) 0.887(0.065) 0.926(0.052) 0.803(0.122) 0.841(0.102)
FPR 0.027(0.006) 0.031(0.008) 0.031(0.008) 0.024(0.006) 0.029(0.009) 0.132(0.020) 0.091(0.021)
S2 with the imbalanced design
CE 0.005(0.009) 0.077(0.140) 0.078(0.140) - - 0.208(0.066) 0.208(0.066)
PME 4.713(0.381) 6.938(5.715) 7.252(6.652) 4.282(0.570) 7.838(0.587) 7.081(1.942) 8.155(1.178)
TPR 0.912(0.045) 0.830(0.165) 0.820(0.163) 0.889(0.061) 0.922(0.056) 0.884(0.104) 0.911(0.067)
FPR 0.025(0.007) 0.031(0.009) 0.030(0.009) 0.023(0.007) 0.028(0.010) 0.120(0.021) 0.086(0.027)
S3 with the imbalanced design
CE 0.013(0.009) 0.158(0.174) 0.162(0.176) - - 0.274(0.120) 0.274(0.120)
PME 4.449(0.381) 7.160(4.339) 7.243(4.615) 4.958(1.616) 8.249(1.091) 6.989(1.190) 7.981(0.816)
TPR 0.932(0.030) 0.775(0.192) 0.765(0.194) 0.817(0.171) 0.886(0.099) 0.877(0.089) 0.920(0.044)
FPR 0.022(0.006) 0.033(0.010) 0.032(0.009) 0.021(0.009) 0.025(0.012) 0.115(0.020) 0.083(0.017)
Table 3: Simulation results for the Erdös-Rényi network under the scenarios with $K=2$. In each cell, mean (SD) based on 100 replicates.
proposed BCGS SCAN TGemBag TJGL FMGemBag FMJGL
S1 with the balanced design
CE 0.005(0.012) 0.027(0.072) 0.027(0.073) - - 0.193(0.071) 0.193(0.071)
PME 3.513(0.245) 3.658(0.861) 3.634(0.857) 2.919(0.388) 5.146(0.685) 7.077(2.002) 6.940(1.593)
TPR 0.962(0.036) 0.938(0.083) 0.935(0.083) 0.954(0.046) 0.947(0.048) 0.769(0.164) 0.766(0.149)
FPR 0.024(0.005) 0.027(0.007) 0.027(0.006) 0.024(0.006) 0.020(0.006) 0.140(0.031) 0.069(0.020)
S2 with the balanced design
CE 0.008(0.012) 0.059(0.108) 0.056(0.100) - - 0.218(0.116) 0.218(0.116)
PME 3.477(0.240) 3.807(1.067) 3.764(1.001) 2.851(0.380) 5.130(0.629) 5.589(1.822) 5.996(1.695)
TPR 0.971(0.032) 0.935(0.103) 0.932(0.097) 0.963(0.043) 0.952(0.046) 0.879(0.149) 0.873(0.136)
FPR 0.028(0.008) 0.030(0.008) 0.029(0.008) 0.026(0.008) 0.023(0.008) 0.132(0.037) 0.067(0.024)
S3 with the balanced design
CE 0.023(0.030) 0.135(0.172) 0.142(0.176) - - 0.251(0.142) 0.251(0.142)
PME 3.199(0.239) 4.276(1.707) 4.373(1.973) 2.768(0.296) 4.956(0.627) 5.157(1.151) 5.497(1.145)
TPR 0.975(0.023) 0.895(0.154) 0.886(0.156) 0.970(0.032) 0.960(0.038) 0.912(0.079) 0.916(0.081)
FPR 0.024(0.008) 0.030(0.009) 0.029(0.008) 0.026(0.008) 0.022(0.009) 0.139(0.033) 0.074(0.031)
S1 with the imbalanced design
CE 0.005(0.008) 0.033(0.070) 0.034(0.071) - - 0.230(0.092) 0.230(0.092)
PME 3.787(0.379) 4.062(0.932) 4.055(0.951) 3.253(0.582) 5.185(0.681) 6.182(1.978) 6.591(2.380)
TPR 0.933(0.051) 0.884(0.108) 0.880(0.107) 0.904(0.072) 0.938(0.050) 0.813(0.199) 0.819(0.191)
FPR 0.027(0.007) 0.028(0.007) 0.027(0.007) 0.025(0.007) 0.024(0.008) 0.130(0.032) 0.074(0.022)
S2 with the imbalanced design
CE 0.009(0.014) 0.064(0.114) 0.057(0.099) - - 0.195(0.098) 0.195(0.098)
PME 3.689(0.337) 4.200(1.240) 4.140(1.131) 3.122(0.517) 5.186(0.655) 6.258(1.715) 6.543(1.687)
TPR 0.951(0.042) 0.893(0.111) 0.888(0.107) 0.925(0.070) 0.947(0.056) 0.829(0.142) 0.838(0.124)
FPR 0.026(0.008) 0.030(0.009) 0.029(0.009) 0.026(0.008) 0.027(0.010) 0.123(0.025) 0.071(0.016)
S3 with the imbalanced design
CE 0.028(0.030) 0.142(0.149) 0.145(0.150) - - 0.265(0.116) 0.265(0.116)
PME 3.416(0.241) 5.082(2.511) 5.170(2.652) 3.072(0.373) 5.007(0.624) 4.896(0.902) 5.317(0.813)
TPR 0.962(0.026) 0.825(0.163) 0.812(0.162) 0.930(0.058) 0.958(0.036) 0.940(0.054) 0.947(0.060)
FPR 0.023(0.007) 0.031(0.009) 0.030(0.009) 0.027(0.008) 0.026(0.010) 0.131(0.033) 0.076(0.028)
As the similarity between the networks increases (from S1 to S3), it is more challenging to identify subgroup memberships. However, the proposed approach still shows more prominent advantages compared to SCAN, BCGS, FMGemBag, and FMJGL in terms of both heterogeneity and network identification performance. It is interesting that, under S3, although with a certain level of clustering errors, the proposed approach has better network identification performance than TGemBag and TJGL with “ideal” true subgroup memberships, suggesting the effectiveness of the proposed similarity-based prior. Under the scenarios with an imbalanced design or a larger number of subgroups, performance of all approaches declines. Under these scenarios, the proposed approach, when compared to TGemBag, exhibits a slightly worse performance in network identification and estimation due to the more complex unknown heterogeneity and the relatively limited sample size. However, it still consistently outperforms the other alternatives. It maintains its superiority under the scenarios with various types of network structure, including the more challenging power-law networks with denser structures and hub structures. With relatively simpler nearest neighbor networks, when all methods generally perform well with higher TPR values, the proposed approach retains significantly fewer FPs compared to most alternatives.
Additionally, to gain a deeper understanding of the similarity-based prior, we illustrate a representative replicate under the scenario with setting S3 and an imbalanced design. The heatmaps depicting the true sparsity structures and the estimated results of the precision matrices with our proposed approach, SCAN, and TGemBag, are presented in Figures S3-S4 (supplementary materials). The findings indicate that our proposed approach excels in accurately identifying true sparsity structures and more effectively accommodating the similarity structures across different networks.
Following previous studies [11, 12, 31], the main simulations assume that the number of subgroups $K$ is known for a more fair comparison. When $K$ is unknown, we recommend using the BIC for model selection. As an illustration, we consider scenarios with a power-law network and balanced design and evaluate the empirical performance of BIC over 100 replicates with candidate values $K=1,\ldots,5$. The selection frequencies, summarized in Table S5 (supplementary materials), indicate that BIC generally performs well. In particular, the S3 configuration, which exhibits greater subgroup similarity than S1 and S2, poses more challenges for accurate identification. For the BIC-selected values of $K$, Table S6 (supplementary material) reports the corresponding identification and estimation results. In cases where $K$ is incorrectly estimated, label alignment is performed minimizing the estimation error of $\bm{\Psi}$ [22]. The results remain consistent with those presented in Table 2 and Table S2 (supplementary materials).
4 Data Analysis
We illustrate the proposed approach in constructing gene networks of different subgroups of cutaneous melanoma (SKCM) from The Cancer Genome Atlas (TCGA) data. Following existing literature [1], we consider the metastatic subjects and use the mRNA gene expression data and overall survival time for the predictors and response, respectively. $360$ subjects are available with $19,039$ gene expression measurements. Among them, $53.6\%$ are non-censored, with the observed survival times ranging from 2.6 to 357.1 months (median 44.05 months). As the sample size is relatively small, we follow the existing studies [12] and conduct gene filtering. Specifically, we first identify 2,260 genes with significance levels below $0.01$ from univariate Cox-regression models and then focus on those present in six pathways (pyrimidine metabolism, P53 signaling, cysteine and methionine metabolism, chemokine signaling, cell cycle, and hedgehog signaling pathways) that have demonstrated important implications for SKCM in the literature. With this preprocessing, 91 genes are obtained for downstream analysis.
As the number of subgroups $K$ is unknown, we consider the candidate set $\left\{1,2,3,4,5,6,7\right\}$ and adopt the BIC criterion for selecting the optimal one. With the proposed approach, the BIC criterion leads to two subgroups with sizes of 128 (subgroup 1) and 232 (subgroup 2). The two networks consist of 181 and 345 edges among 80 and 83 genes, respectively. Among them, 52 edges are common and all have the same signs. A graphical representation is provided in Figure 1, and more detailed estimation results are provided in the supplementary materials. To gain deeper insights into the identified networks, we conduct Gene Ontology (GO) enrichment analysis using DAVID 2021 to examine the functional and biological connections of the related genes. Some significantly enriched GO terms are obtained for the common edges, and genes that exhibit unique edges in the two subgroups are found to be linked to different GO terms. The sensible biological findings provide support to the validity of the proposed network estimation. We refer to Section S7 of the supplementary materials for the detailed GO enrichment analysis results.
Figure 1: Data analysis: gene networks for the two subgroups identified by the proposed approach. In each network, the black edges are shared by the two subgroups, and the colored edges are unique in the corresponding subgroup.
Comparative analysis using the alternative approaches is also carried out with $K=2$. The summary results are reported in Table S7 (supplementary materials), which provide the number of subjects and network edges in the two subgroups identified by different approaches and their overlaps. The subgroups identified using different approaches are matched by correlation. In Table S7 (supplementary materials), different approaches are observed to identify subgroups with different subjects. To get a deeper understanding of the heterogeneity analysis performance, we examine whether these subgroups are associated with survival time. In Figure 2, we present the Kaplan–Meier (KM) curves of survival time of different subgroups identified by the proposed and alternative approaches. It is observed that the KM curves of the two subgroups are significantly distinguished from each other for the proposed approach (P value: 0.005). Comparatively, the FM approach shows less significant differences. BCGS and SCAN approaches fail to identify subgroups with distinct survival. These biologically reasonable findings indicate that the proposed approach can satisfactorily identify survival-related subgroups and also provide support for its effectiveness in clinical subgroup identification.
Figure 2: Data analysis: Kaplan-Meier curves for the two subgroups identified by different approaches.
In Table S7 (supplementary materials), different approaches demonstrate a moderate number of overlapping edges. To provide indirect support for the network estimation results, we use the resampling strategy and randomly divide the data into a training and a testing set. As in the literature [9], the values of the negative log-likelihood statistic (NLS) for the 91 genes are calculated to evaluate prediction accuracy. The average NLS values of 100 resamplings are 7102.118 for the proposed approach, 7658.492 for BCGS, 7674.739 for SCAN, 8044.482 for FMGemBag, and 8221.576 for FMJGL, which suggest a satisfactory network estimation accuracy of the proposed approach.
5 Discussion
In this article, we have introduced a novel Bayesian approach for jointly learning multiple sparse networks in a supervised paradigm when the subgroup memberships are unknown. Benefiting from the Bayesian framework, the proposed approach can effectively accommodate both the heterogeneous distributions of predictors and the heterogeneous associations between predictors and survival time, leading to results with stronger clinical implications. In addition to the sparse priors for network estimation, we have introduced a similarity prior to encourage similar network sparsity as well as similar positive or negative effects across different subgroups, which can more effectively address the similarity of subgroups. We have thoroughly examined the theoretical properties of the proposed approach. A series of numerical experiments have been conducted to demonstrate the superior performance of the proposed approach in both heterogeneity and network identification. By applying the approach to TCGA data, we have discovered various gene networks for SKCM patients and identified biologically relevant relationships associated with clinically sensible subgroups.
In this study, we have employed the commonly used Gaussian mixture models for heterogeneity analysis, which show satisfactory performance. It can be of interest to explore alternative robust techniques, such as the exponential family graphical models proposed by [27] or other nonparametric graphical models, for studying heavy-tailed data. For survival modeling, we employ the Normal-error AFT model as a canonical example. Our framework is general and can be extended to more flexible survival models, such as the Weibull AFT model by appropriately modifying the likelihood function, representing a valuable direction for future research. Our focus in this data analysis has been mainly on the identification of gene networks using expression data. There is vast potential for a more comprehensive understanding of cancer mechanisms by incorporating other omics measurements, such as mutation and DNA methylation.
6 Disclosure statement
The authors have the following conflicts of interest to declare.
7 Data Availability Statement
The data that support the findings in this paper are openly available in TCGA (The Cancer Genome Atlas) at https://portal.gdc.cancer.gov/projects/TCGA-SKCM.
SUPPLEMENTARY MATERIAL
Supplement.pdf
Supplement to “Supervised Bayesian Joint Graphical Model for Simultaneous Network Estimation and Subgroup Identification”, including a graphical illustration of the method, algorithm details, sensitivity results, definition of $W$ in Assumption 2, technical proofs, simulation settings and results, and additional real data results.
superviseNetcode.zip
The R package superviseNet implementing the proposed method.
Realdata.xlsx
Network estimation results for real data using the proposed approach.